2. Company A purchases $200, 000 of equipment in year 0. It decides to use straight line depreciation over the expected 20 ye show yoir work
Answer #1
Equipment Cost $                    2,00,000.00
Useful life 20 Years
Straight Line Depreciation=(Cost-Salvage Value) /Useful life
Depreciation=($200000-0)/20= $                        10,000.00
Annual Depreciation tax shield=($10000*40%)=(A) $                          4,000.00
P.V of annuity 20 years @ 14%=(B) 6.62313
Present worh of the after tax Depreciation recovery($4000*6.62313)=(A)*(B) $                        26,492.52
Ans (C ) Present worh of the after tax Depreciation recovery $                        26,500.00 nearest dollar